OpenStudy (anonymous):

Determine if triangle XYZ with coordinates X (1, 1), Y (5, 6), and Z (6, 2) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\huge\color{blue}{ d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} }\) use this distance formula to find all the sides first.

OpenStudy (solomonzelman):

After you find them, plug them into \(\huge\color{red}{a^2+b^2=c^2 }\) the biggest side will be what you plug in for c. Now, GO FOR IT !

OpenStudy (anonymous):

can you draw it out for me please

OpenStudy (solomonzelman):

No need, just find the distances.

OpenStudy (anonymous):

i drew it already

OpenStudy (anonymous):

ok im about to

OpenStudy (anonymous):

so i plug all of the coordinates into th distance formula correct

OpenStudy (anonymous):

wait that makes no sense what would x2 and y2 be we only have 3 sets of coordinates